OpenStudy (anonymous):

there is an 80% chance that a person eats dinner a 70% chance a person eats dessert and a 40% chance the person will eat dinner and dessert. which of the following is true

OpenStudy (kropot72):

@ElToro48 You need to post the choices if you expect somebody to help you.

OpenStudy (anonymous):

A.eating dinner and eating dessert are dependent events because p(dinner) -p(dessert) =0.8-0.7=0.1 which is less than p(dinner and dessert) = 0.4 B.Eating dinner and eating dessert are dependent events because p(dinner)x p(dessert)= 0.8 x 0.7=0.56 which is not equal to p(dinner and dessert) =0.4 C.eating dinner and eating dessert are independent events because p(dinner) x p(dessert)=0.8 x 0.7=0.56 which is not equal to p(dinner and dessert) = 0.4 D.eating dinner and eating dessert are independent events because p(dinner) - p(dessert)=0.8 - 0.7=0.1 which is less than p(dinner and dessert) = 0.4

OpenStudy (anonymous):

@kropot72^^^

OpenStudy (kropot72):

Independent events have no effect on the occurrence (or non-occurrence) of each other. Events A and B are independent if, and only if \[\large P(A\cap B)=P(A) \times P(B)\] You should now be able to choose correctly.